OpenStudy (anonymous):

solve the exponent tial equation express the solution set in terms of natural logarithims

OpenStudy (anonymous):

4^x+4=5^2x+5

OpenStudy (anonymous):

ln4^x+4=ln5^2x+5

OpenStudy (anonymous):

x+4ln4=2x+5ln5

OpenStudy (anonymous):

here I am a bit confused

OpenStudy (alekos):

Don't you mean 4^(x+4)=5^(2x+5) ?